1.The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.
2.A bicyclist rode into the country for 5 h. In returning, her
speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
3. A bus leaves a station at 1 P.M., traveling west at an average
rate of 44 mi/h. One hour later a second bus leaves the same station, traveling
east at a rate of 48 mi/h. At what time will the two buses be 274 mi apart?
4. Mika leaves Boston for Baltimore at 10:00 A.M., traveling at
45 mi/h. One hour later, Hiroko leaves Baltimore for Boston on the same route, traveling
at 50 mi/h. If the two cities are 425 mi apart, when will Mika and Hiroko meet?

Answer by Paul(988)

1. 2(2x-5+x)=230
6x-10=230
6x=240
x=40
2(40)-5=75
Hence, the width is 40cm and the length is 75cm.

2.5x=4(x+5)
5x=4x+20
x=20
20+5 = 25
Hence her two speeds were 20miles per hours and 25 miles per hour.

3.44(t+1)+48t=274
44t+44+48t=274
92t=230
t=2.5
Hence, at 2.5h they will be 274 miles apart.

4.45(t+1)+50t=425
45t+45+50t=425
95t=380
t=4
10:00+4 = 2:00p.m.
Hence after 4h or at 2:00 p.m. ; they will meet.